A room contains air at 27C and 101 kPa at a relative humidity of 88 percent. determine, using four decimals:

(a) the partial pressure of dry air. Pa (b) the humidity ratio, W.
р, о p = Pa + pv 0.622 рр. т. Pete а at T = 27°С, Рg = 3.567 kPa PV = kPa а) Pa = kPa b) =
Answer #1

$=0.88 d = Pro Prs 0.88= Pr. 3.567 Pr= 3.13 90 kPa 0 W = 0. 622 Pr P-Pv P= PatPg Pa= = 101 KPA P= 10 + 3.567= 104.567 kPa 0.6